This is the question: A plane flies 720 mi against a steady 30-mi/h headwind
and then returns to the same point with the wind.
If the entire trip takes 10 hrs., what is the plane's speed in the air?
Please help. Thank you.

Answer by checkley75(3666)

720/(S-30)+720/(S+30)=10
[720(S+30)+720(S-30)]/(S-30)(S+30)=10
(720S+21600+720S-21600)/(S^2-900)=10 COMBINE TERMS & CROSS MULTIPLY
1440S=10S^2-9000
10S^2-1440S-9000=0
S^2-144S-900=0 NOW FACTOR:
(X-150)(X+6)=0
X-150=0
X=150 ANSWER FOR THE SPEED OF THE PLANE IN STILL AIR.
PROOF
720/120+720/180=10
6+4=10
10=10